Aluminum will react with oxygen to form aluminum oxide. It takes 2.00 L of pure oxygen gas at STP to react completely with a certain sample of aluminum. What is the mass of aluminum reacted?

 


1
Expert's answer
2021-04-19T04:39:27-0400

Find moles of O2 and then from the balanced equation find moles and mass of Al.

PV = nRT

"n =\frac{ PV}{RT} =\frac{ (1 atm)(2.00 L)}{(0.0821 Latm/Kmol)(273 K)}"


"n = moles ...O_2 = 0.0892 moles"


"moles Al = 0.0892 moles (O_2)×\frac{ 4 moles( Al)}{3 }"


moles O2 = 0.119 moles Al


mass Al = 0.119 moles x 26.98 g/mole = 3.21 g of Al